### v1

# There may be a unique $Q$-point

Abstract. We show that in the model obtained by iteratively pseudo-intersecting a Ramsey ultrafilter via a length-$\omega_{2}$ countable support iteration of restricted Mathias forcing over a ground model satisfying CH, there is a unique $Q$-point up to isomorphism.
key-words: $Q$-point, Ramsey ultrafilter, Mathias forcing
2010 Mathematics Subject Classification: 03E35 03E17
1 Introduction
Throughout this paper, read ultrafilter as non-principal ultrafilter on $\omega$. For $x\subseteq\omega$, we denote by $[x]^{\omega}$ the set of infinite subsets of $x$ and by $\operatorname{fin}(x)$ the set of finite subsets of $x$.
Recall that an ultrafilter $E$ is a $Q$-point if and only if for every interval partition $\{[k_{i},k_{i+1}):i\in\omega\}$ of $\omega$, there exists some $x\in E$ such that $\forall i\in\omega:|x\cap[k_{i},k_{i+1})|\leq 1$. Furthermore, an ultrafilter $\mathcal{U}$ is a Ramsey ultrafilter if and only if the Maiden has no winning strategy in the ultrafilter game for $\mathcal{U}$, played between the Maiden and Death:
Definition 1.1.
Let $\mathcal{U}$ be an ultrafilter. The ultrafilter game for $\mathcal{U}$ proceeds as follows:
The Maiden opens the game and plays some $y_{0}\in\mathcal{U}$. Death responds by playing some $n_{0}\in y_{0}$. In the $(k+1)$-th move, the Maiden having played $y_{0}\supseteq y_{1}\supseteq...\supseteq y_{k}$, and Death having played $n_{0}<n_{1}<...<n_{k}$, the Maiden plays some $y_{k+1}\in[y_{k}]^{\omega}\cap\mathcal{U}$, and Death responds by playing some $n_{k+1}\in y_{k+1}$, $n_{k+1}>n_{k}$.
Death wins if and only if $\{n_{i}:i\in\omega\}\in\mathcal{U}$.
[…]
Definition 1.2.
Let $\mathcal{U}$ be a Ramsey ultrafilter. Mathias forcing restricted to $\mathcal{U}$, written $\mathbb{M}_{\mathcal{U}}$, consists of conditions $\langle s,x\rangle\in\text{fin}(\omega)\times\mathcal{U}$ with $\max s<\min x$, ordered by
$\langle s,x\rangle\leq_{\mathbb{M}_{\mathcal{U}}}\langle t,y\rangle:\iff s\subseteq t\land x\supseteq y\land t\setminus s\subseteq x.$
Note that we use the convention that stronger forcing conditions are larger. The forcing notion $\mathbb{M}_{\mathcal{U}}$ clearly satisfies the c.c.c. and is therefore proper. We will need the following additional facts.
Fact 1.3 (e.g., see [1, Theorem 26.3]).
Let $\mathcal{U}$ be a Ramsey ultrafilter. The forcing notion $\mathbb{M}_{\mathcal{U}}$ has the pure decision property, i.e., for any sentence $\varphi$ in the forcing language and any $\mathbb{M}_{\mathcal{U}}$-condition $\langle s,x\rangle$, there exists $y\in[x]^{\omega}\cap\mathcal{U}$ such that either $\langle s,y\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}}\varphi$ or $\langle s,y\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}}\neg\varphi$.
Definition 1.4.
Recall that a forcing notion $\mathbb{P}$ has the Laver property iff for every $\mathbb{P}$-name $\undertilde{g}$ for an element of ${{}^{\omega}}\omega$ such that there exists $f\in{{}^{\omega}}\omega\cap\mathbf{V}$ with
$\mathbb{P}\Vdash\forall n\in\omega:\undertilde{g}(n)\leq f(n),$
we have that $\mathbb{P}$ forces that there exists $c:\omega\to\text{fin}(\omega)$ in $\mathbf{V}$ with
$\forall n\in\omega:|c(n)|\leq 2^{n}\text{ and }\undertilde{g}(n)\in c(n).$
Fact 1.5 (e.g., see [1, Corollary 26.8]).
Let $\mathcal{U}$ be a Ramsey ultrafilter. The forcing notion $\mathbb{M}_{\mathcal{U}}$ has the Laver property.
Fact 1.6 (e.g., see [2, Ch. VI, 2.10D]).
The Laver property is preserved under countable support iterations of proper forcing notions.
2 Result
Main Theorem.
It is consistent that there exists a unique $Q$-point, and this $Q$-point is a Ramsey ultrafilter.
Proof.
Assume that the ground model $\mathbf{V}$ satisfies CH. By induction, we define:
  • (i)
    A countable support iteration $\mathbb{P}_{\omega_{2}}:=\langle\mathbb{P}_{\xi},,\undertilde{Q}_{\xi}:\xi\in\omega_{2}\rangle$ of c.c.c. forcing notions,
  • (ii)
    A sequence $\langle\undertilde{\mathcal{U}}_{\xi}:\xi\in\omega_{2}\rangle$, such that
    $\forall\xi\in\omega_{2}:\mathbb{P}_{\xi}\Vdash``\undertilde{\mathcal{U}}_{\xi}\text{ is a Ramsey ultrafilter extending }\bigcup_{\iota\in\xi}\undertilde{\mathcal{U}}_{\iota}"$
    and $\undertilde{Q}_{\xi}$ is a $\mathbb{P}_{\xi}$-name for Mathias forcing restricted to $\undertilde{\mathcal{U}}_{\xi}$,
Assume that we are in step $\xi\in\omega_{2}$. Let $G_{\xi}$ be $\mathbb{P}_{\xi}$-generic over $\mathbf{V}$ and work in $\mathbf{V}[G_{\xi}]$. Note that since $\mathbb{P}_{\xi}$ is a countable support iteration of proper forcing notions that are forced to be of size $\leq\omega_{1}$, we have $\mathbf{V}[G_{\xi}]\models\textsf{CH}$ (e.g., see [3, Theorem 2.12]). For each $\iota\in\xi$, let $\eta_{\iota}$ be the Mathias real added at stage $\iota$.
If $\xi=\xi^{\prime}+1$, $\eta_{\xi^{\prime}}$ pseudo-intersects $\undertilde{\mathcal{U}}_{\xi^{\prime}}[G_{\xi}]$ and we may construct a Ramsey ultrafilter on $\eta_{\xi^{\prime}}$ using CH (and extend it to $\omega$ to obtain $\mathcal{U}_{\xi}$). Similarly, if $\xi$ is a limit ordinal and $\text{cf}(\xi)=\omega$, we can build $\mathcal{U}_{\xi}$ on a pseudo-intersection of the tower $\langle\eta_{\iota}:\iota\in\xi\rangle$. Finally, if $\text{cf}(\xi)=\omega_{1}$, then $\bigcup_{\iota\in\xi}\undertilde{\mathcal{U}}_{\iota}[G_{\xi}]$ is already a Ramsey ultrafilter, since no new reals are added at stage $\xi$. For the same reason we also have that $\mathcal{U}_{\omega_{2}}:=\bigcup_{\xi\in\omega_{2}}\undertilde{\mathcal{U}}_{\xi}[G]$ is a Ramsey ultrafilter in $\mathbf{V}[G]$, where $G$ is $\mathbb{P}_{\omega_{2}}$-generic over $\mathbf{V}$.
Fact 2.1 (e.g., see [3, Theorem 2.10]).
$\mathbb{P}_{\omega_{2}}$ is proper and satisfies the $\omega_{2}$-c.c..
We need to show that $\mathcal{U}_{\omega_{2}}$ is the only $Q$-point in $\mathbf{V}[G]$. To see this, assume by contradiction that $\mathbf{V}[G]\models``E\text{ is a }Q\text{-point and not isomorphic to }\mathcal{U}_{\omega_{2}}"$.
Lemma 2.2.
There exists $\delta\in\omega_{2}$ such that $E\cap\mathbf{V}[G_{\delta}]\in\mathbf{V}[G_{\delta}]$ and $\mathbf{V}[G_{\delta}]\models``E\cap\mathbf{V}[G_{\delta}]\text{ is a }Q\text{-point and not isomorphic to }\mathcal{U}_{\delta}"$.
Proof.
Fix $\xi\in\omega_{2}$ and consider names $\undertilde{e}_{\xi}$, $\undertilde{i}_{\xi}$, $\undertilde{s}_{\xi}$, $\undertilde{b}_{\xi}$ and $\undertilde{f}_{\xi}$ such that $\mathbb{P}_{\omega_{2}}$ forces that
  • (i)
    ``$\undertilde{e}_{\xi}$ is an enumeration (in $\omega_{1}$) of $\undertilde{E}\cap\mathbf{V}[\underaccent{\dot}{G}_{\xi}]"$. For each $\alpha\in\omega_{1}$ and $n\in\omega$ let $\mathcal{E}_{\xi,\alpha,n}\subseteq\mathbb{P}_{\omega_{2}}$ be a maximal antichain deciding ``$n\in\undertilde{e}_{\xi}(\alpha)"$.
  • (ii)
    ``$\undertilde{i}_{\xi}$ is an enumeration (in $\omega_{1}$) of the set of interval partitions of $\omega$ in $\mathbf{V}[\underaccent{\dot}{G}_{\xi}]"$. Note that we may assume that $\undertilde{i}_{\xi}$ is a $\mathbb{P}_{\xi}$-name.
  • (iii)
    “For all $\alpha\in\omega_{1}$, $\undertilde{s}_{\xi}(\alpha)$ is an element of $\undertilde{E}$ that intersects each interval in the interval partition $\undertilde{i}_{\xi}(\alpha)$ in at most one point”. Let $\mathcal{S}_{\xi,\alpha,n}\subseteq\mathbb{P}_{\omega_{2}}$ be a maximal antichain deciding ``$n\in\undertilde{s}_{\xi}(\alpha)"$.
  • (iv)
    ``$\undertilde{b}_{\xi}$ is an enumeration (in $\omega_{1}$) of all permutations of $\omega$ in $\mathbf{V}[\underaccent{\dot}{G}_{\xi}]"$. We may again assume that $\undertilde{b}_{\xi}$ is a $\mathbb{P}_{\xi}$-name.
  • (v)
    “For all $\alpha\in\omega_{1}$, $\undertilde{f}_{\xi}(\alpha)$ is a pair $\text{op}(\undertilde{x}_{\alpha},\undertilde{y}_{\alpha})$ such that $\undertilde{x}_{\alpha}$ is in $\undertilde{E}$, $\undertilde{y}_{\alpha}$ is in $\undertilde{\mathcal{U}}_{\omega_{2}}$ and $\undertilde{b}_{\xi}(\alpha)[\undertilde{x}_{\alpha}]$ is disjoint from $\undertilde{y}_{\alpha}"$. Let $\mathcal{X}_{\xi,\alpha,n}\subseteq\mathbb{P}_{\omega_{2}}$ be a maximal antichain deciding ``$n\in\undertilde{x}_{\alpha}"$, and define $\mathcal{Y}_{\xi,\alpha,n}$ analogously.
By the $\omega_{2}$-c.c. of $\mathbb{P}_{\omega_{2}}$, there exists for each $\xi\in\omega_{2}$ some $\gamma_{\xi}\in\omega_{2}$ greater than $\xi$ such that all the above antichains consist of $\mathbb{P}_{\gamma_{\xi}}$-conditions. Recursively define $\lambda(0)=0$, $\lambda(\xi+1)=\gamma_{\lambda(\xi)}$ and for limit ordinals $\xi:\lambda(\xi)=\bigcup_{\iota\in\xi}\lambda(\iota)$, for $\xi\leq\omega_{1}$. Set $\delta:=\lambda(\omega_{1})$ and consider the extension $\mathbf{V}[G_{\delta}]$. Since $\text{cf}(\delta)=\omega_{1}$, we have that $E\cap\mathbf{V}[G_{\delta}]=\bigcup_{\iota\in\omega_{1}}E\cap\mathbf{V}[G_{\lambda(\iota)}]$, and since each $E\cap\mathbf{V}[G_{\lambda(\iota)}]$ is an element of $\mathbf{V}[G_{\delta}]$ by (i), $E\cap\mathbf{V}[G_{\delta}]$ is an element of $\mathbf{V}[G_{\delta}]$ (and an ultrafilter). Furthermore, any interval partition of $\omega$ in $\mathbf{V}[G_{\delta}]$ already appears in some $\mathbf{V}[G_{\lambda(\iota)}]$, $\iota\in\omega_{1}$, where it equals $\undertilde{i}_{\lambda(\iota)}[G_{\lambda(\iota)}](\alpha)$ for some $\alpha\in\omega_{1}$. Since $\undertilde{s}_{\lambda(\iota)}[G_{\delta}](\alpha)\in E\cap\mathbf{V}[G_{\delta}]$, we obtain that $E\cap\mathbf{V}[G_{\delta}]$ is a $Q$-point. Finally and analogously, any permutation of $\omega$ in $\mathbf{V}[G_{\delta}]$ already appears in $\mathbf{V}[G_{\lambda(\iota)}]$ for some $\iota\in\omega_{1}$ and hence there are witnesses $\undertilde{x}_{\alpha}[G_{\delta}]\in E\cap\mathbf{V}[G_{\delta}]$ and $\undertilde{y}_{\alpha}[G_{\delta}]\in\mathcal{U}_{\omega_{2}}\cap\mathbf{V}[G_{\delta}]=\mathcal{U}_{\delta}$ witnessing that $E\cap\mathbf{V}[G_{\delta}]$ and $\mathcal{U}_{\delta}$ are not isomorphic.   $\dashv$
We now designate $\mathbf{V}[G_{\delta}]$ as the new ground model and rename the $Q$-point $E\cap\mathbf{V}[G_{\delta}]$ to $E$ and the Ramsey ultrafilter $\mathcal{U}_{\delta}$ to $\mathcal{U}$. Note that by the Factor-Lemma (e.g., see [4, Theorem 4.6]), the quotient $\mathbb{P}_{\omega_{2}}/G_{\delta}$ is again isomorphic to a countable support iteration of restricted Mathias forcings. In particular, by Facts 1.5 and 1.6, $\mathbb{P}_{\omega_{2}}/G_{\delta}$ is isomorphic to the two-step iteration $\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}$, where $\mathbb{M}_{\mathcal{U}}\Vdash``\undertilde{R}\text{ has the Laver property}"$.
It remains to show the following.
Proposition 2.3.
Let $E$ be a $Q$-point and $\mathcal{U}$ a Ramsey ultrafilter such that $E$ and $\mathcal{U}$ are not isomorphic. Let $\mathbb{M}_{\mathcal{U}}$ be Mathias forcing restricted to $\mathcal{U}$ and let $\undertilde{R}$ be a $\mathbb{M}_{\mathcal{U}}$-name such that $\mathbb{M}_{\mathcal{U}}\Vdash``\undertilde{R}$ has the Laver property”. Then $\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}\Vdash``E$ cannot be extended to a $Q$-point”.
Proof.
It suffices to show that if $\langle p,\undertilde{q}\rangle\in\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}$ and a $\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}$-name $\undertilde{a}$ for a strictly increasing element of ${{}^{\omega}}\omega$ are such that
$\langle p,\undertilde{q}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}}\forall n\in\omega:\undertilde{a}(n)\in(\underaccent{\dot}{\eta}(n-1),\underaccent{\dot}{\eta}(n)],$
then there exists some $v\in E$ and some $\langle\bar{p},\undertilde{\bar{q}}\rangle$ greater than $\langle p,\undertilde{q}\rangle$ such that
$\langle\bar{p},\undertilde{\bar{q}}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}}|\text{range}(\undertilde{a})\cap v|<\omega.$
Recall that $\underaccent{\dot}{\eta}$ is the canonical $\mathbb{M}_{\mathcal{U}}$-name for the Mathias real (assume $\mathbb{M}_{\mathcal{U}}\Vdash\underaccent{\dot}{\eta}(-1)=-\infty$).
Note that $\undertilde{a}$ is forced by $\mathbb{M}_{\mathcal{U}}$ to be dominated by $\underaccent{\dot}{\eta}$. Hence, by the Laver property of $\undertilde{R}$, there exists a $\mathbb{M}_{\mathcal{U}}$-name $\undertilde{c}$ for a function from $\omega$ to $\text{fin}(\omega)$ and some $\langle p^{\prime},\undertilde{q}^{\prime}\rangle\geq_{\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}}\langle p,\undertilde{q}\rangle$ such that
$\langle p^{\prime},\undertilde{q}^{\prime}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}}\forall n\in\omega:\undertilde{a}(n)\in\undertilde{c}(n)\text{ and }|\undertilde{c}(n)|\leq 2^{n}.$
We may assume without loss of generality that $p^{\prime}\Vdash_{\mathbb{M}_{\mathcal{U}}}\forall n\in\omega:\undertilde{c}(n)\subseteq(\underaccent{\dot}{\eta}(n-1),\underaccent{\dot}{\eta}(n)]$. Let $\undertilde{C}$ be a $\mathbb{M}_{\mathcal{U}}$-name for an element of $[\omega]^{\omega}$ such that $p^{\prime}\Vdash_{\mathbb{M}_{\mathcal{U}}}\undertilde{C}=\bigcup\text{range}(\undertilde{c})$. Hence we have
$\langle p^{\prime},\undertilde{q}^{\prime}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}}\forall n\in\omega:\undertilde{a}(n)\in\undertilde{C}\cap(\underaccent{\dot}{\eta}(n-1),\underaccent{\dot}{\eta}(n)]\text{ and }|\undertilde{C}\cap(\underaccent{\dot}{\eta}(n-1),\underaccent{\dot}{\eta}(n)]|\leq 2^{n}.$
Lemma 2.4.
Write $p^{\prime}=\langle s,x_{0}\rangle$. There exists $x_{1}\in[x_{0}]^{\omega}\cap\mathcal{U}$ such that the $\mathbb{M}_{\mathcal{U}}$-condition $\langle s,x_{1}\rangle\geq_{\mathbb{M}_{\mathcal{U}}}\langle s,x_{0}\rangle$ has the following property:
For every $t\in\text{fin}(x_{1})$, there exists $C_{t}\in\text{fin}(\omega)$ such that
$\langle s\cup t,x_{1}\setminus(\max t)^{+}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}}\undertilde{C}\cap(\max t)^{+}=C_{t}.$
Proof.
We define a strategy for the Maiden in the ultrafilter game for $\mathcal{U}$, which will not be a winning strategy since $\mathcal{U}$ is a Ramsey ultrafilter.
Since $\mathbb{M}_{\mathcal{U}}$ has pure decision, there exists $C_{\emptyset}\subseteq(\max s)^{+}$ and $y_{0}\in[x_{0}]^{\omega}\cap\mathcal{U}$ such that $\langle s,y_{0}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}}\undertilde{C}\cap(\max s)^{+}=C_{\emptyset}$. The Maiden starts by playing $y_{0}$.
Assume $y_{0}\supseteq y_{1}\supseteq...\supseteq y_{k}$ and $n_{0}<n_{1}<...<n_{k}$ have been played, where $\forall i\leq k:y_{i}\in\mathcal{U}$ and $n_{i}\in y_{i}$. Again by pure decision, for each $t\subseteq\{n_{0},n_{1},...,n_{k}\}$ with $\max t=n_{k}$, there exists $z_{t}\in[y_{k}\setminus n_{k}^{+}]^{\omega}\cap\mathcal{U}$ and $C_{t}\subseteq n_{k}^{+}$ such that $\langle s\cup t,z_{t}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}}\undertilde{C}\cap(n_{k})^{+}=C_{t}$. The Maiden plays
$y_{k+1}:=\bigcap_{\begin{subarray}{c}{t\subseteq\{n_{i}:i\leq k\}}\\ {\max t=n_{k}}\end{subarray}}z_{t}.$
Since Death wins, we have that $x_{1}:=\{n_{i}:i\in\omega\}\in\mathcal{U}$. It is easy to check that this $x_{1}$ satisfies the lemma.   $\dashv$
The following lemma strengthens the previous one.
Lemma 2.5.
Assume $\langle s,x_{1}\rangle$ is as in the conclusion of the previous lemma. There exists $x_{2}\in[x_{1}]^{\omega}\cap\mathcal{U}$ such that $\langle s,x_{2}\rangle$ has the following property:
For every $t\in\text{fin}(x_{2})$, every $m\in x_{2}\setminus\max t$ and all $n,n^{\prime}\in x_{2}\setminus m^{+}$, it holds that $C_{t\cup\{n\}}\cap m^{+}=C_{t\cup\{n^{\prime}\}}\cap m^{+}$.
Proof.
We again prove this by playing the ultrafilter game for $\mathcal{U}$. Assume $y_{0}:=x_{1}\supseteq y_{1}\supseteq...\supseteq y_{k}$ and $n_{0}<n_{1}<...<n_{k}$ have been played. For every $t\subseteq\{n_{0},n_{1},...,n_{k}\}$ and every $d\subseteq n_{k}^{+}$ consider the set
$P_{t,d}:=\{n\in y_{k}\setminus n_{k}^{+}:C_{t\cup\{n\}}\cap n_{k}^{+}=d\}.$
Note that for every $t\subseteq\{n_{0},n_{1},...,n_{k}\}$, the set $\{P_{t,d}:d\subseteq n_{k}^{+}\}$ is a partition of $y_{k}\setminus n_{k}^{+}$ into finitely many pieces. Hence, there exists one $d_{t}\subseteq n_{k}^{+}$ such that $P_{t,d_{t}}\in\mathcal{U}$.
The Maiden plays
$y_{k+1}:=\bigcap_{t\subseteq\{n_{i}:i\leq k\}}P_{t,d_{t}}.$
Death will win and hence $x_{2}:=\{n_{i}:i\in\omega\}\in\mathcal{U}$. It is again not hard to check that $x_{2}$ satisfies the lemma.   $\dashv$
The following fact will be needed later.
Fact 2.6.
Without loss of generality, we may assume that for all $n\in\{\max s\}\cup x_{2}$, if $n$ is the $j$’th element of $s\cup x_{2}$ in increasing order, then $n>2^{j+1}$.
Proof.
Note that the conclusion of Lemmas 2.4 and 2.5 also holds for each $\langle s^{\prime},x^{\prime}\rangle\geq_{\mathbb{M}_{\mathcal{U}}}\langle s,x_{2}\rangle$. Hence we simply trim $x_{2}$ such that the enumeration of $s\cup x_{2}$ dominates $2^{j+1}$ above $|s|$ and replace $s$ with $s\cup\{\min x_{2}\}$ and $x_{2}$ with $x_{2}\setminus\{\min x_{2}\}$.   $\dashv$
Next, let $N$ be a countable elementary submodel of some large enough $\mathcal{H}_{\chi}$ such that $\{\mathcal{U},\mathbb{M}_{\mathcal{U}},\undertilde{C},\langle s,x_{2}\rangle\}\in N$. By induction, construct a sequence $N_{0}\subseteq N_{1}\subseteq...$ of finite subsets of $N$ such that
  • (i)
    $\{\mathcal{U},\mathbb{M}_{\mathcal{U}},\undertilde{C},\langle s,x_{2}\rangle,s,x_{2}\}\subseteq N_{0}$,
  • (ii)
    $\bigcup_{i\in\omega}N_{i}=N$,
  • (iii)
    $\forall i\in\omega:k_{i}:=N_{i}\cap\omega\in\omega$.
  • (iv)
    $\forall i\in\omega:\forall t\in\text{fin}(\omega):t\in N_{i}\iff t\subseteq N_{i}$,
  • (v)
    If $\langle m,l,D\rangle\in(\omega\times\omega\times\text{fin}(\omega))\cap N_{i}$, then $m,l,D\in N_{i}$ (and hence $D\subseteq N_{i}$ by the previous condition).
  • (vi)
    $\forall i\in\omega:$ If $\varphi(x,a_{0},...,a_{l})$ is a formula of length less than $2025$ with $a_{0},...,a_{l}\in N_{i}$ and $N\models\exists x\varphi(x,a_{0},...,a_{l})$, then there exists $b\in N_{i+1}$ such that $N\models\varphi(b,a_{0},...,a_{l})$.
Lemma 2.7.
$\langle s,x_{2}\rangle$ forces that
$\forall i\in\omega\setminus\{0,1\}:\undertilde{C}\setminus(\max s)^{+}\cap[k_{i-1},k_{i})\neq\emptyset\implies\begin{cases}\text{range}(\underaccent{\dot}{\eta})\cap[k_{i-2},k_{i-1})\neq\emptyset,\text{ or}\\ \text{range}(\underaccent{\dot}{\eta})\cap[k_{i-1},k_{i})\neq\emptyset,\text{ or}\\ \text{range}(\underaccent{\dot}{\eta})\cap[k_{i},k_{i+1})\neq\emptyset.\end{cases}$
Proof.
Assume $\langle s\cup t,x^{\prime}\rangle\geq_{\mathbb{M}_{\mathcal{U}}}\langle s,x_{2}\rangle$, $a\in\omega\setminus(\max s)^{+}$ and $i\in\omega\setminus\{0,1\}$ are such that
$\langle s\cup t,x^{\prime}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}}a\in\undertilde{C}\setminus(\max s)^{+}\cap[k_{i-1},k_{i}).$
We show that $\langle s\cup t,x^{\prime}\rangle$ forces one of the three possible conclusions in the statement of the lemma.
By possibly extending $t$, we may assume that $t$ contains at least one element that is greater than $a$. Let $l_{0}:=\max(t\cap a)$ and $l^{*}:=\min(t\setminus a)$. Furthermore, let $m^{*}:=\max(x_{2}\cap l^{*})$. Hence, $l_{0}$ and $l^{*}$ are consecutive elements of $t$ and $l_{0}\leq m^{*}<l^{*}$ and $l_{0}<a\leq l^{*}$. We distinguish between two cases:
Case I.
Assume $l_{0}\leq m^{*}\leq a\leq l^{*}$.
If $l^{*}\in[k_{i-1},k_{i})$, we are done, since this means that $\langle s\cup t,x^{\prime}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}}l^{*}\in\text{range}(\underaccent{\dot}{\eta})\cap[k_{i-1},k_{i})$. Hence, assume $l^{*}\notin[k_{i-1},k_{i})$, which means that $l^{*}\notin N_{i}$, since $l^{*}$ is certainly not in $N_{i-1}$ (if it were, $a$ would be as well by (iii)). Note that $l^{*}$ witnesses that
$N\models\exists l:l=\min(x_{2}\setminus a).$
Hence, by (v), we have that $l^{*}\in N_{i+1}$ and thus $l^{*}\in[k_{i},k_{i+1})$.
Case II.
Assume $l_{0}<a<m^{*}<l^{*}$.
Let $t^{\prime}:=t\cap a$, i.e., $l_{0}:=\max t^{\prime}$, and let $i^{*}\in\omega\setminus\{0\}$ be such that $l_{0}\in[k_{i^{*}-1},k_{i^{*}})$, i.e., $l_{0}$ first appears in $N_{i^{*}}$. If $i^{*}=i$, we are again done, hence assume that $a\notin N_{i^{*}}$. We will show that $i^{*}=i-1$.
Let $j\in\omega$ be such that $l^{*}$ is the $j$’th elements of $s\cup t$ in increasing order. By Lemmas 2.4 and 2.5, there is $C_{t^{\prime}\cup\{l^{*}\}}\subseteq(l^{*})^{+}$ such that
$\langle s\cup t^{\prime}\cup\{l^{*}\},x_{2}\setminus(l^{*})^{+}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}}\undertilde{C}\cap(l^{*})^{+}=C_{t^{\prime}\cup\{l^{*}\}}.$
Set $D^{*}:=C_{t^{\prime}\cup\{l^{*}\}}\cap(l_{0},m^{*})$. Since
$\langle s\cup t^{\prime}\cup\{l^{*}\},x_{2}\setminus(l^{*})^{+}\rangle\leq_{\mathbb{M}_{\mathcal{U}}}\langle s\cup t,x^{\prime}\rangle,$
and since $l_{0}<a<m^{*}$ by assumption, we must have $a\in D^{*}$. Furthermore, note that $D^{*}\subseteq C_{t^{\prime}\cup\{l^{*}\}}\cap(l_{0},l^{*}]$ and thus $|D^{*}|=:\gamma\leq 2^{j}$.
Now, $m^{*}$, $l^{*}$ and $D^{*}$ witness that
$N\models\exists\langle m,l,D\rangle:\begin{cases}m,l\in x_{2}\setminus{l_{0}}^{+},m<l,\text{ and}\\ D\subseteq(l_{0},m),\text{ and }\\ |D|=\gamma,\text{ and }\\ \langle s\cup t^{\prime}\cup\{l\},x_{2}\setminus l^{+}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}}\undertilde{C}\cap(l_{0},m)=D.\end{cases}$
Since $l_{0}$ is the $(j-1)$’th element of $s\cup t^{\prime}$, we have $l_{0}>2^{j}$ by Fact 2.6.333Note that the additional requirement in Fact 2.6 that $\max s$ is already larger than $2^{|s|}$ is needed here, since $l_{0}$ could be $\max s$. Hence, since $l_{0}\in N_{i^{*}}$, it follows that $\gamma\in N_{i^{*}}$. Thus, all the parameters in the above formula lie in $N_{i^{*}}$, which implies that there exists $\langle m^{\dagger},l^{\dagger},D^{\dagger}\rangle\in N_{i^{*}+1}$ satisfying the formula.
Claim. $l^{\dagger}\geq a$
Note that the proof of this claim will finish the proof of the Lemma, since $l^{\dagger}\in N_{i^{*}+1}$ by (v) and thus $a\in N_{i^{*}+1}\setminus N_{i^{*}}$.
Proof. Assume by contradiction that $l^{\dagger}<a$, i.e.,
$l_{0}<m^{\dagger}<l^{\dagger}<a<m^{*}<l^{*}.$
By Lemma 2.5, we have that
$C_{t^{\prime}\cup\{l^{\dagger}\}}\cap(m^{\dagger})=C_{t^{\prime}\cup\{l^{*}\}}\cap(m^{\dagger}).$
Since $\langle s\cup t^{\prime}\cup\{l^{\dagger}\},x_{2}\setminus(l^{\dagger})^{+}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}}\undertilde{C}\cap(l_{0},m^{\dagger})=D^{\dagger}$, it follows that $C_{t^{\prime}\cup\{l^{*}\}}\cap(m^{\dagger})=D^{\dagger}$ and hence $D^{\dagger}=D^{*}\cap(l_{0},m^{\dagger})$. However, both $D^{\dagger}$ and $D^{*}$ have size $\gamma$ and thus $D^{*}\subseteq(l_{0},m^{\dagger})$, which is a contradiction to the fact that $a\in D^{*}$ and $a>m^{\dagger}$. $\blacksquare$   $\dashv$
We now only need one final lemma to finish the proof of the proposition and thus of the main theorem.
Lemma 2.8.
Let $I:=\{[k_{i},k_{i+1}):i\in\omega\}$ be any interval partition of $\omega$ and $E$ and $\mathcal{U}$ non-isomorphic $Q$-points. Then there exist $v\in E$ and $u\in\mathcal{U}$ such that
$\forall i\in\omega\setminus\{0\}:v\cap[k_{i},k_{i+1})\neq\emptyset\implies\begin{cases}u\cap[k_{i-1},k_{i})=\emptyset,\text{ and}\\ u\cap[k_{i},k_{i+1})=\emptyset,\text{ and}\\ u\cap[k_{i+1},k_{i+2})=\emptyset.\end{cases}$
Proof.
Say that a $Q$-point element selects from an interval partition if it intersects each interval in exactly one point. Let $v_{0}\in E$ and $u_{0}\in\mathcal{U}$ be such that they select from $I$. Let $f$ be an order-preserving bijection from $v_{0}$ to $u_{0}$, extended to a permutation of $\omega$. Thus, for each $i\in\omega$, $f$ sends the element selected by $v_{0}$ in $[k_{i},k_{i+1})$ to the element selected by $u_{0}$ in $[k_{i},k_{i+1})$. Since $E$ and $\mathcal{U}$ are non-isomorphic, there exist $v_{1}\in[v_{0}]^{\omega}\cap E$ and $u_{1}\in[u_{0}]^{\omega}\cap\mathcal{U}$ such that $u_{1}\cap f[v_{1}]=\emptyset$. Hence, for all $i\in\omega\setminus\{0\}$:
$v_{1}\cap[k_{i},k_{i+1})\neq\emptyset\implies u_{1}\cap[k_{i},k_{i+1})=\emptyset.$
Both $E$ and $\mathcal{U}$ contain the set
$y_{\varepsilon}:=\bigcup_{\begin{subarray}{c}{i\in\omega}\\ {i\equiv\varepsilon\;(\text{mod}\;3)}\end{subarray}}[k_{i},k_{i+1}),$
each for exactly one $\varepsilon=\varepsilon(E),\varepsilon({\mathcal{U}})\in 3$. Let $v_{2}:=v_{1}\cap y_{\varepsilon(E)}\in E$ and $u_{2}:=u_{1}\cap y_{\varepsilon(\mathcal{U})}\in\mathcal{U}$. If $\varepsilon(E)=\varepsilon(\mathcal{U})$ then $v_{2}$ and $u_{2}$ satisfy the lemma, hence assume without loss of generality that $\varepsilon(E)=0$ and $\varepsilon(\mathcal{U})=1$.
Let $\bar{v}_{0}\in E$ and $\bar{u}_{0}\in\mathcal{U}$ be elements that select from the interval partition
$\{[k_{i},k_{i+2}):i\in\omega,i\equiv 0\;(\text{mod}\;3)\}\cup\{[k_{i},k_{i+1}):i\in\omega,i\equiv 2\;(\text{mod}\;3)\}.$
Again, by considering a permutation of $\omega$ that maps the element selected by $\bar{v}_{0}$ in any interval to the element selected by $\bar{u}_{0}$ in the same interval, we find $\bar{v}_{1}\in[\bar{v}_{0}]^{\omega}\cap E$ and $\bar{u}_{1}\in[\bar{u}_{0}]^{\omega}\cap\mathcal{U}$ such that $\bar{v}_{1}$ and $\bar{u}_{1}$ never select from the same interval. Now, clearly, $v_{1}\cap\bar{v}_{1}\in E$ and $u_{1}\cap\bar{u}_{1}\in\mathcal{U}$ work.   $\dashv$
We can now finish the proof of the proposition and hence of the main theorem: Let $v\in E$, $u\in\mathcal{U}$ be given by the previous lemma for the interval partition $\{[k_{i},k_{i+1}):i\in\omega\}\cup\{[0,k_{0})\}$ constructed in the proof of Lemma 2.7. Let $G\ast H$ be any $\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}$-generic filter containing $\langle\langle s,x_{2}\rangle,\undertilde{q}^{\prime}\rangle$. By Lemma 2.7, we have that in $\mathbf{V}[G\ast H]$, whenever $\text{range}(\undertilde{a}[G\ast H])\setminus(\max s)^{+}$ intersects one of the intervals $[k_{i},k_{i+1})$, then the Mathias real $\eta$ intersects $[k_{i},k_{i+1})$ or one of the adjacent intervals $[k_{i-1},k_{i})$ or $[k_{i+1},k_{i+2})$. Since $\text{range}(\eta)$ is almost contained in $u$, the same is true for $u$ in place of $\eta$ above some $n\geq(\max s)^{+}$. Hence, $\text{range}(\undertilde{a}[G\ast H])\setminus n$ is disjoint from $v$.   $\dashv$   $\dashv$

### v2

@@
-There may be a unique $Q$-point
+A unique $Q$-point and infinitely many near-coherence classes of ultrafilters
 Abstract.
 We show that in the model obtained by iteratively pseudo-intersecting a Ramsey ultrafilter via a length-$\omega_{2}$ countable support iteration of restricted Mathias forcing over a ground model satisfying CH, there is a unique $Q$-point up to isomorphism.
+In particular, it is consistent that there is only one $Q$-point while there are $2^{\mathfrak{c}}$-many near-coherence classes of ultrafilters.
 key-words: $Q$-point, Ramsey ultrafilter, Mathias forcing
@@
 […]
+While the consistency of the non-existence of $Q$-points is a well-established fact with a variety of witnesses apart from the Mathias model333such as the Laver and Miller models (see [7] and [11], respectively)., the construction of models containing only ‘few’ $Q$-points seems to have received less attention.
+However, such models do arise naturally as models containing only few near-coherence classes of ultrafilters [^4] Two ultrafilters $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$ are nearly-coherent iff there is some finite-to-one $f\in{{}^{\omega}}{\omega}$ such that $f(\mathcal{U}_{1})=f(\mathcal{U}_{2})$, where $f(\mathcal{U}_{i}):=\{X\subseteq\omega:f^{-1}[X]\in\mathcal{U}_{i}\}$.
+Note that two $Q$-points are nearly-coherent iff they are isomorphic.: Indeed, Mildenberger [10] has constructed models with exactly two and exactly three near-coherence classes, and it is easy to see that these contain exactly one and exactly two $Q$-points, respectively: In her model with exactly two near-coherence classes, one class contains a Ramsey ultrafilter, while the other class contains an ultrafilter that is $\omega_{1}$-generated.
+Hence, this latter class cannot contain a $Q$-point, since her models satisfy $\mathfrak{d}=\mathfrak{c}=\omega_{2}$ and such a $Q$-point would thus have to be ${<}\mathfrak{d}$-generated, which is impossible.
+Analogously, in Mildenberger’s model with exactly three near-coherence classes, two classes are represented by Ramsey ultrafilters, while the third contains an $\omega_{1}$-generated ultrafilter – giving exactly two $Q$-points in total.
+The construction of models with exactly $n$ near-coherence classes of ultrafilters for various finite $n\geq 4$ would similarly yield the consistency of exactly $m$ $Q$-points for some $m<n$ [^5] The inequality is strict since such a model must satisfy $\mathfrak{u}<\mathfrak{d}$, a result due to Banakh and Blass [9].
+Hence, one of the $n$ near-coherence classes contains a ${<}\mathfrak{d}$-generated ultrafilter and thus no $Q$-point.
+The model considered in this paper is of a different nature, however: It contains only one $Q$-point while its number of near-coherence classes is $2^{\mathfrak{c}}$, i.e., the model’s lack of $Q$-points is not the consequence of a lack of near-coherence classes.
+This follows from the fact that dominating reals are added at each of the $\omega_{2}$-many stages of the iteration, which gives $\mathfrak{b}=\mathfrak{d}=\mathfrak{c}=\omega_{2}$ in the final extension.
+Since $\mathfrak{b}\leq\mathfrak{u}$ (see Solomon [8]), we have $\mathfrak{u}=\mathfrak{d}=\omega_{2}$ in our model, and hence there are $2^{\mathfrak{c}}$-many near-coherence classes of ultrafilters by Banakh and Blass [9].
 Definition 1.2.
@@
 Main Theorem.
-It is consistent that there exists a unique $Q$-point, and this $Q$-point is a Ramsey ultrafilter.
+It is consistent that there is a unique $Q$-point while there are $2^{\mathfrak{c}}$-many near-coherence classes of ultrafilters.
 Proof.
@@
 Let $\undertilde{C}$ be a $\mathbb{M}_{\mathcal{U}}$-name for an element of $[\omega]^{\omega}$ such that $p^{\prime}\Vdash_{\mathbb{M}_{\mathcal{U}}}\undertilde{C}=\bigcup\text{range}(\undertilde{c})$.
-Hence we have
+Hence, we have
 $\langle p^{\prime},\undertilde{q}^{\prime}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}}\forall n\in\omega:\undertilde{a}(n)\in\undertilde{C}\cap(\underaccent{\dot}{\eta}(n-1),\underaccent{\dot}{\eta}(n)]\text{ and }|\undertilde{C}\cap(\underaccent{\dot}{\eta}(n-1),\underaccent{\dot}{\eta}(n)]|\leq 2^{n}.$
@@
 Note that the conclusion of Lemmas 2.4 and 2.5 also holds for each $\langle s^{\prime},x^{\prime}\rangle\geq_{\mathbb{M}_{\mathcal{U}}}\langle s,x_{2}\rangle$.
-Hence we simply trim $x_{2}$ such that the enumeration of $s\cup x_{2}$ dominates $2^{j+1}$ above $|s|$ and replace $s$ with $s\cup\{\min x_{2}\}$ and $x_{2}$ with $x_{2}\setminus\{\min x_{2}\}$.
+Hence, we simply trim $x_{2}$ such that the enumeration of $s\cup x_{2}$ dominates $2^{j+1}$ above $|s|$ and replace $s$ with $s\cup\{\min x_{2}\}$ and $x_{2}$ with $x_{2}\setminus\{\min x_{2}\}$.
 $\dashv$
@@
 $N\models\exists\langle m,l,D\rangle:\begin{cases}m,l\in x_{2}\setminus{l_{0}}^{+},m<l,\text{ and}\\ D\subseteq(l_{0},m),\text{ and }\\ |D|=\gamma,\text{ and }\\ \langle s\cup t^{\prime}\cup\{l\},x_{2}\setminus l^{+}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}}\undertilde{C}\cap(l_{0},m)=D.\end{cases}$
-Since $l_{0}$ is the $(j-1)$’th element of $s\cup t^{\prime}$, we have $l_{0}>2^{j}$ by Fact 2.6.333Note that the additional requirement in Fact 2.6 that $\max s$ is already larger than $2^{|s|}$ is needed here, since $l_{0}$ could be $\max s$.
+Since $l_{0}$ is the $(j-1)$’th element of $s\cup t^{\prime}$, we have $l_{0}>2^{j}$ by Fact 2.6.666Note that the additional requirement in Fact 2.6 that $\max s$ is already larger than $2^{|s|}$ is needed here, since $l_{0}$ could be $\max s$.
 Hence, since $l_{0}\in N_{i^{*}}$, it follows that $\gamma\in N_{i^{*}}$.

### v3

@@
-A unique $Q$-point and infinitely many near-coherence classes of ultrafilters
+arXiv:2505.17960v3[math.LO] 01 Mar 2026
 Abstract.
@@
 Throughout this paper, read ultrafilter as non-principal ultrafilter on $\omega$.
-For $x\subseteq\omega$, we denote by $[x]^{\omega}$ the set of infinite subsets of $x$ and by $\operatorname{fin}(x)$ the set of finite subsets of $x$.
+For $x\subseteq\omega$, we denote by $x]^{\omega}$ the set of infinite subsets of $x$ and by $x]^{<\omega}$ the set of finite subsets of $x$.
 Recall that an ultrafilter $E$ is a $Q$-point if and only if for every interval partition $\{[k_{i},k_{i+1}):i\in\omega\}$ of $\omega$, there exists some $x\in E$ such that $\forall i\in\omega:|x\cap[k_{i},k_{i+1})|\leq 1$.
@@
 Death wins if and only if $\{n_{i}:i\in\omega\}\in\mathcal{U}$.
-[…]
+It is well-known that every Ramsey ultrafilter is a $Q$-point.
+Canjar [4] showed that the existence of $2^{\mathfrak{c}}$-many Ramsey ultrafilters follows from the assumption $\mathit{cov}(\mathcal{M})=\mathfrak{c}$.
+The weaker assumption $\mathit{cov}(\mathcal{M})=\mathfrak{d}$ implies the existence of $2^{\mathfrak{c}}$ $Q$-points, as was shown by Millán [12].
+It is well-known that in the Mathias model – the model obtained by a length-$\omega_{2}$ countable support iteration of unrestricted Mathias forcing over a ground model satisfying CH – there are no $Q$-points (see [9, Proposition 26.23]).
+In fact, the Mathias model contains no rapid ultrafilters, where an ultrafilter $E$ is rapid if for every $f\in{}^{\omega}\omega$ there exists some $x\in E$ such that $\forall n\in\omega:|x\cap f(n)|\leq n$ (note that every $Q$-point is rapid).
+It follows that both the Mathias model and the model considered in this paper satisfy $\mathit{cov}(\mathcal{M})=\omega_{1}<\mathfrak{d}=\mathfrak{c}=\omega_{2}$.
 […]
 While the consistency of the non-existence of $Q$-points is a well-established fact with a variety of witnesses apart from the Mathias model333such as the Laver and Miller models (see [13] and [3], respectively)., the construction of models containing only ‘few’ $Q$-points seems to have received less attention.
-However, such models do arise naturally as models containing only few near-coherence classes of ultrafilters [^4] Two ultrafilters $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$ are nearly-coherent iff there is some finite-to-one $f\in{{}^{\omega}}{\omega}$ such that $f(\mathcal{U}_{1})=f(\mathcal{U}_{2})$, where $f(\mathcal{U}_{i}):=\{X\subseteq\omega:f^{-1}[X]\in\mathcal{U}_{i}\}$.
-Note that two $Q$-points are nearly-coherent iff they are isomorphic.: Indeed, Mildenberger [10] has constructed models with exactly two and exactly three near-coherence classes, and it is easy to see that these contain exactly one and exactly two $Q$-points, respectively: In her model with exactly two near-coherence classes, one class contains a Ramsey ultrafilter, while the other class contains an ultrafilter that is $\omega_{1}$-generated.
-Hence, this latter class cannot contain a $Q$-point, since her models satisfy $\mathfrak{d}=\mathfrak{c}=\omega_{2}$ and such a $Q$-point would thus have to be ${<}\mathfrak{d}$-generated, which is impossible.
+However, such models do arise naturally as models containing only few near-coherence classes of ultrafilters [^4] Two ultrafilters $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$ are nearly-coherent if there is some finite-to-one $f\in{{}^{\omega}}{\omega}$ such that $f(\mathcal{U}_{1})=f(\mathcal{U}_{2})$, where $f(\mathcal{U}_{i}):=\{X\subseteq\omega:f^{-1}[X]\in\mathcal{U}_{i}\}$.
+Note that two $Q$-points are nearly-coherent if and only if they are isomorphic.: Indeed, Mildenberger [11] has constructed models with exactly two and exactly three near-coherence classes, and it is easy to see that these contain exactly one and exactly two $Q$-points, respectively: In her model with exactly two near-coherence classes, one class contains a Ramsey ultrafilter, while the other class contains an ultrafilter that is $\omega_{1}$-generated.
+Hence, this latter class cannot contain a $Q$-point, since her models satisfy $\mathfrak{d}=\mathfrak{c}=\omega_{2}$ and such a $Q$-point would thus have to be $\mathfrak{d}-$generated, which is impossible.
 Analogously, in Mildenberger’s model with exactly three near-coherence classes, two classes are represented by Ramsey ultrafilters, while the third contains an $\omega_{1}$-generated ultrafilter – giving exactly two $Q$-points in total.
-The construction of models with exactly $n$ near-coherence classes of ultrafilters for various finite $n\geq 4$ would similarly yield the consistency of exactly $m$ $Q$-points for some $m<n$ [^5] The inequality is strict since such a model must satisfy $\mathfrak{u}<\mathfrak{d}$, a result due to Banakh and Blass [9].
-Hence, one of the $n$ near-coherence classes contains a ${<}\mathfrak{d}$-generated ultrafilter and thus no $Q$-point.
+The construction of models with exactly $n$ near-coherence classes of ultrafilters for various finite $n\geq 4$ would similarly yield the consistency of exactly $m$ $Q$-points for some m<nm<n.
+[^5] The inequality is strict since such a model must satisfy $\mathfrak{u}<\mathfrak{d}$, a result due to Banakh and Blass [2].
+Hence, one of the $n$ near-coherence classes contains a $\mathfrak{d}-$generated ultrafilter and thus no $Q$-point.
+[^6] See the note on the next page.
 The model considered in this paper is of a different nature, however: It contains only one $Q$-point while its number of near-coherence classes is $2^{\mathfrak{c}}$, i.e., the model’s lack of $Q$-points is not the consequence of a lack of near-coherence classes.
-This follows from the fact that dominating reals are added at each of the $\omega_{2}$-many stages of the iteration, which gives $\mathfrak{b}=\mathfrak{d}=\mathfrak{c}=\omega_{2}$ in the final extension.
+This follows from the fact that dominating reals are added at each of the $\omega_{2}$ stages of the iteration, which gives $\mathfrak{b}=\mathfrak{d}=\mathfrak{c}=\omega_{2}$ in the final extension.
 Since $\mathfrak{b}\leq\mathfrak{u}$ (see Solomon [15]), we have $\mathfrak{u}=\mathfrak{d}=\omega_{2}$ in our model, and hence there are $2^{\mathfrak{c}}$-many near-coherence classes of ultrafilters by Banakh and Blass [2].
+Incidentally, both our model as well as Mildenberger’s models answer the two questions posed in a recent paper of Raghavan [5].
+He asked whether it is consistent that there are $Q$-points, while (1) there is no Tukey maximal $Q$-point, and (2) the $Q$-points are not cofinal in the RK-ordering.
+[^7] See, for example, [6] and [7] on the Tukey- and RK-orderings of ultrafilters.
+Note that the unique $Q$-point in our model is simultaneously a $P$-point, and these are never Tukey maximal (see [7, Corollary 19]).
+Note added in proof:
+Based on the technique used below, in the forthcoming paper [10], Halbeisen, Horvath, and Özalp have established the consistency of the statement “There are exactly $n$ $Q$-points up to isomorphism and $\mathfrak{u}=\mathfrak{d}=\mathfrak{c}$, for any $n\in\omega$.
+1 Definitions and Preliminaries
+Before we state and prove our main result in the next section, let us give some definitions, relations, and related results.
 Definition 1.1.
 Let $\mathcal{U}$ be a Ramsey ultrafilter.
-Mathias forcing restricted to $\mathcal{U}$, written $\mathbb{M}_{\mathcal{U}}$, consists of conditions $\langle s,x\rangle\in\text{fin}(\omega)\times\mathcal{U}$ with $\max s<\min x$, ordered by
-$\langle s,x\rangle\leq_{\mathbb{M}_{\mathcal{U}}}\langle t,y\rangle:\iff s\subseteq t\land x\supseteq y\land t\setminus s\subseteq x.$
-Note that we use the convention that stronger forcing conditions are larger.
-The forcing notion $\mathbb{M}_{\mathcal{U}}$ clearly satisfies the c.c.c.
-and is therefore proper.
-We will need the following additional facts.
-Fact 1.3 (e.g., see [1, Theorem 26.3]).
+Mathias forcing restricted to $\mathcal{U}$, written $\mathbb{M}_{\mathcal{U}}$, consists of conditions $\langle s,x\rangle\in[\omega]^{<\omega}\times\mathcal{U}$ with $\max s<\min x$, ordered by
+$\langle s,x\rangle\leq_{\mathbb{M}_{\mathcal{U}}}\langle t,y\rangle:\iff s\supseteq t\land x\subseteq y\land s\setminus t\subseteq y.$
+The forcing notion $\mathbb{M}_{\mathcal{U}}$ clearly satisfies the c.c.c. and is therefore proper. We will need the following additional facts.
+Fact 1.2 (e.g., see [9, Theorem 26.3]).
 Let $\mathcal{U}$ be a Ramsey ultrafilter.
@@
 Definition 1.3.
-Recall that a forcing notion $\mathbb{P}$ has the Laver property iff for every $\mathbb{P}$-name $\undertilde{g}$ for an element of ${{}^{\omega}}\omega$ such that there exists $f\in{{}^{\omega}}\omega\cap\mathbf{V}$ with
+Recall that a forcing notion $\mathbb{P}$ has the Laver property if for every $\mathbb{P}$-name $\undertilde{g}$ for an element of ${{}^{\omega}}\omega$ such that there exists $f\in{{}^{\omega}}\omega\cap\mathbf{V}$ with
 $\mathbb{P}\Vdash\forall n\in\omega:\undertilde{g}(n)\leq f(n),$
-we have that $\mathbb{P}$ forces that there exists $c:\omega\to\text{fin}(\omega)$ in $\mathbf{V}$ with
+we have that $\mathbb{P}$ forces that there exists $c:\omega\to[\omega]^{<\omega}$ in $\mathbf{V}$ with
 $\forall n\in\omega:|c(n)|\leq 2^{n}\text{ and }\undertilde{g}(n)\in c(n).$
-Fact 1.5 (e.g., see [1, Corollary 26.8]).
+Fact 1.4 (e.g., see [9, Corollary 26.8]).
 Let $\mathcal{U}$ be a Ramsey ultrafilter. The forcing notion $\mathbb{M}_{\mathcal{U}}$ has the Laver property.
-Fact 1.6 (e.g., see [2, Ch. VI, 2.10D]).
+Fact 1.5 (e.g., see [14, Ch. VI, 2.10D]).
 The Laver property is preserved under countable support iterations of proper forcing notions.
 2 Result
+Now, we are ready to state the
 Main Theorem.
 It is consistent that there is a unique $Q$-point while there are $2^{\mathfrak{c}}$-many near-coherence classes of ultrafilters.
-Proof.
-Assume that the ground model $\mathbf{V}$ satisfies CH. By induction, we define:
+The proof is given by the following construction and the subsequent results: Assume that the ground model $\mathbf{V}$ satisfies CH. By induction, we define:
   • (i)
-    A countable support iteration $\mathbb{P}_{\omega_{2}}:=\langle\mathbb{P}_{\xi},,\undertilde{Q}_{\xi}:\xi\in\omega_{2}\rangle$ of c.c.c. forcing notions,
+    A countable support iteration $\mathbb{P}_{\omega_{2}}:=\langle\mathbb{P}_{\xi},\undertilde{Q}_{\xi}:\xi\in\omega_{2}\rangle$ of c.c.c. forcing notions,
   • (ii)
@@
 Let $G_{\xi}$ be $\mathbb{P}_{\xi}$-generic over $\mathbf{V}$ and work in $\mathbf{V}[G_{\xi}]$.
-Note that since $\mathbb{P}_{\xi}$ is a countable support iteration of proper forcing notions that are forced to be of size $\leq\omega_{1}$, we have $\mathbf{V}[G_{\xi}]\models\textsf{CH}$ (e.g., see [3, Theorem 2.12]).
+Note that since $\mathbb{P}_{\xi}$ is a countable support iteration of proper forcing notions that are forced to be of size $\leq\omega_{1}$, we have $\mathbf{V}[G_{\xi}]\models\textsf{CH}$ (e.g., see [1, Theorem 2.12]).
 For each $\iota\in\xi$, let $\eta_{\iota}$ be the Mathias real added at stage $\iota$.
@@
 For the same reason we also have that $\mathcal{U}_{\omega_{2}}:=\bigcup_{\xi\in\omega_{2}}\undertilde{\mathcal{U}}_{\xi}[G]$ is a Ramsey ultrafilter in $\mathbf{V}[G]$, where $G$ is $\mathbb{P}_{\omega_{2}}$-generic over $\mathbf{V}$.
-Fact 2.1 (e.g., see  [3, Theorem 2.10]).
-$\mathbb{P}_{\omega_{2}}$ is proper and satisfies the $\omega_{2}$-c.c..
-We need to show that $\mathcal{U}_{\omega_{2}}$ is the only $Q$-point in $\mathbf{V}[G]$.
-To see this, assume by contradiction that $\mathbf{V}[G]\models``E\text{ is a }Q\text{-point and not isomorphic to }\mathcal{U}_{\omega_{2}}"$.
+Fact 2.1 (see, e.g., [1, Theorem 2.10]).
+$\mathbb{P}_{\omega_{2}}$ is proper and satisfies $\omega_{2}$-$c.$.
+[…]
 Lemma 2.2.
-There exists $\delta\in\omega_{2}$ such that $E\cap\mathbf{V}[G_{\delta}]\in\mathbf{V}[G_{\delta}]$ and $\mathbf{V}[G_{\delta}]\models``E\cap\mathbf{V}[G_{\delta}]\text{ is a }Q\text{-point and not isomorphic to }\mathcal{U}_{\delta}"$.
+[…]
 Proof.
@@
   • (i)
-    ``$\undertilde{e}_{\xi}$ is an enumeration (in $\omega_{1}$) of $\undertilde{E}\cap\mathbf{V}[\underaccent{\dot}{G}_{\xi}]"$.
-For each $\alpha\in\omega_{1}$ and $n\in\omega$ let $\mathcal{E}_{\xi,\alpha,n}\subseteq\mathbb{P}_{\omega_{2}}$ be a maximal antichain deciding ``$n\in\undertilde{e}_{\xi}(\alpha)"$.
+    $\undertilde{e}_{\xi}$ is an enumeration (in $\omega_{1}$) of $\undertilde{E}\cap\mathbf{V}[\underaccent{\dot}{G}_{\xi}]$".
+For each $\alpha\in\omega_{1}$ and $n\in\omega$ let $\mathcal{E}_{\xi,\alpha,n}\subseteq\mathbb{P}_{\omega_{2}}$ be a maximal antichain deciding $n\in\undertilde{e}_{\xi}(\alpha)$".
   • (ii)
-    ``$\undertilde{i}_{\xi}$ is an enumeration (in $\omega_{1}$) of the set of interval partitions of $\omega$ in $\mathbf{V}[\underaccent{\dot}{G}_{\xi}]"$.
+    $\undertilde{i}_{\xi}$ is an enumeration (in $\omega_{1}$) of the set of interval partitions of $\omega$ in $\mathbf{V}[\underaccent{\dot}{G}_{\xi}]$".
 Note that we may assume that $\undertilde{i}_{\xi}$ is a $\mathbb{P}_{\xi}$-name.
@@
     “For all $\alpha\in\omega_{1}$, $\undertilde{s}_{\xi}(\alpha)$ is an element of $\undertilde{E}$ that intersects each interval in the interval partition $\undertilde{i}_{\xi}(\alpha)$ in at most one point”.
-Let $\mathcal{S}_{\xi,\alpha,n}\subseteq\mathbb{P}_{\omega_{2}}$ be a maximal antichain deciding ``$n\in\undertilde{s}_{\xi}(\alpha)"$.
+Let $\mathcal{S}_{\xi,\alpha,n}\subseteq\mathbb{P}_{\omega_{2}}$ be a maximal antichain deciding $n\in\undertilde{s}_{\xi}(\alpha)$".
   • (iv)
-    ``$\undertilde{b}_{\xi}$ is an enumeration (in $\omega_{1}$) of all permutations of $\omega$ in $\mathbf{V}[\underaccent{\dot}{G}_{\xi}]"$.
+    $\undertilde{b}_{\xi}$ is an enumeration (in $\omega_{1}$) of all permutations of $\omega$ in $\mathbf{V}[\underaccent{\dot}{G}_{\xi}]$".
 We may again assume that $\undertilde{b}_{\xi}$ is a $\mathbb{P}_{\xi}$-name.
   • (v)
-    “For all $\alpha\in\omega_{1}$, $\undertilde{f}_{\xi}(\alpha)$ is a pair $\text{op}(\undertilde{x}_{\alpha},\undertilde{y}_{\alpha})$ such that $\undertilde{x}_{\alpha}$ is in $\undertilde{E}$, $\undertilde{y}_{\alpha}$ is in $\undertilde{\mathcal{U}}_{\omega_{2}}$ and $\undertilde{b}_{\xi}(\alpha)[\undertilde{x}_{\alpha}]$ is disjoint from $\undertilde{y}_{\alpha}"$.
-Let $\mathcal{X}_{\xi,\alpha,n}\subseteq\mathbb{P}_{\omega_{2}}$ be a maximal antichain deciding ``$n\in\undertilde{x}_{\alpha}"$, and define $\mathcal{Y}_{\xi,\alpha,n}$ analogously.
-By the $\omega_{2}$-c.c.
-of $\mathbb{P}_{\omega_{2}}$, there exists for each $\xi\in\omega_{2}$ some $\gamma_{\xi}\in\omega_{2}$ greater than $\xi$ such that all the above antichains consist of $\mathbb{P}_{\gamma_{\xi}}$-conditions.
+    “For all $\alpha\in\omega_{1}$, $\undertilde{f}_{\xi}(\alpha)$ is a pair $\langle\undertilde{x}_{\alpha},\undertilde{y}_{\alpha}\rangle$ such that $\undertilde{x}_{\alpha}$ is in $\undertilde{E}$, $\undertilde{y}_{\alpha}$ is in $\undertilde{\mathcal{U}}_{\omega_{2}}$ and $\undertilde{b}_{\xi}(\alpha)[\undertilde{x}_{\alpha}]$ is disjoint from $\undertilde{y}_{\alpha}$".
+Let $\mathcal{X}_{\xi,\alpha,n}\subseteq\mathbb{P}_{\omega_{2}}$ be a maximal antichain deciding $n\in\undertilde{x}_{\alpha}$", and define $\mathcal{Y}_{\xi,\alpha,n}$ analogously.
+Since $\mathbb{P}_{\omega_{2}}$ satisfies $\omega_{2}$-c.c., there exists for each $\xi\in\omega_{2}$ some $\gamma_{\xi}\in\omega_{2}$ greater than $\xi$ such that all the above antichains consist of $\mathbb{P}_{\gamma_{\xi}}$-conditions.
 Recursively define $\lambda(0)=0$, $\lambda(\xi+1)=\gamma_{\lambda(\xi)}$ and for limit ordinals $\xi:\lambda(\xi)=\bigcup_{\iota\in\xi}\lambda(\iota)$, for $\xi\leq\omega_{1}$.
@@
 We now designate $\mathbf{V}[G_{\delta}]$ as the new ground model and rename the $Q$-point $E\cap\mathbf{V}[G_{\delta}]$ to $E$ and the Ramsey ultrafilter $\mathcal{U}_{\delta}$ to $\mathcal{U}$.
-Note that by the Factor-Lemma (e.g., see [4, Theorem 4.6]), the quotient $\mathbb{P}_{\omega_{2}}/G_{\delta}$ is again isomorphic to a countable support iteration of restricted Mathias forcings.
-In particular, by Facts 1.5 and 1.6, $\mathbb{P}_{\omega_{2}}/G_{\delta}$ is isomorphic to the two-step iteration $\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}$, where $\mathbb{M}_{\mathcal{U}}\Vdash``\undertilde{R}\text{ has the Laver property}"$.
-It remains to show the following.
+Note that by the Factor-Lemma (e.g., see [8, Theorem 4.6]), the quotient $\mathbb{P}_{\omega_{2}}/G_{\delta}$ is again isomorphic to a countable support iteration of restricted Mathias forcings.
+In particular, by Facts 1.4 and 1.5, $\mathbb{P}_{\omega_{2}}/G_{\delta}$ is isomorphic to the two-step iteration $\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}$, where $\mathbb{M}_{\mathcal{U}}\Vdash``\undertilde{R}\text{ has the Laver property}$".
+It remains to show the following
 Proposition 2.3.
@@
 $\langle p,\undertilde{q}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}}\forall n\in\omega:\undertilde{a}(n)\in(\underaccent{\dot}{\eta}(n-1),\underaccent{\dot}{\eta}(n)],$
-then there exists some $v\in E$ and some $\langle\bar{p},\undertilde{\bar{q}}\rangle$ greater than $\langle p,\undertilde{q}\rangle$ such that
+then there exists some $v\in E$ and some $\langle\bar{p},\undertilde{\bar{q}}\rangle\leq_{\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}}\langle p,\undertilde{q}\rangle$ such that
 $\langle\bar{p},\undertilde{\bar{q}}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}}|\text{range}(\undertilde{a})\cap v|<\omega.$
@@
 Note that $\undertilde{a}$ is forced by $\mathbb{M}_{\mathcal{U}}$ to be dominated by $\underaccent{\dot}{\eta}$.
-Hence, by the Laver property of $\undertilde{R}$, there exists a $\mathbb{M}_{\mathcal{U}}$-name $\undertilde{c}$ for a function from $\omega$ to $\text{fin}(\omega)$ and some $\langle p^{\prime},\undertilde{q}^{\prime}\rangle\geq_{\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}}\langle p,\undertilde{q}\rangle$ such that
+Hence, by the Laver property of $\undertilde{R}$, there exists a $\mathbb{M}_{\mathcal{U}}$-name $\undertilde{c}$ for a function from $\omega$ to $\omega]^{<\omega}$ and some $\langle p^{\prime},\undertilde{q}^{\prime}\rangle\leq_{\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}}\langle p,\undertilde{q}\rangle$ such that
 $\langle p^{\prime},\undertilde{q}^{\prime}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}}\forall n\in\omega:\undertilde{a}(n)\in\undertilde{c}(n)\text{ and }|\undertilde{c}(n)|\leq 2^{n}.$
-We may assume without loss of generality that $p^{\prime}\Vdash_{\mathbb{M}_{\mathcal{U}}}\forall n\in\omega:\undertilde{c}(n)\subseteq(\underaccent{\dot}{\eta}(n-1),\underaccent{\dot}{\eta}(n)]$.
-Let $\undertilde{C}$ be a $\mathbb{M}_{\mathcal{U}}$-name for an element of $[\omega]^{\omega}$ such that $p^{\prime}\Vdash_{\mathbb{M}_{\mathcal{U}}}\undertilde{C}=\bigcup\text{range}(\undertilde{c})$.
-Hence, we have
+[…]
 $\langle p^{\prime},\undertilde{q}^{\prime}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}}\forall n\in\omega:\undertilde{a}(n)\in\undertilde{C}\cap(\underaccent{\dot}{\eta}(n-1),\underaccent{\dot}{\eta}(n)]\text{ and }|\undertilde{C}\cap(\underaccent{\dot}{\eta}(n-1),\underaccent{\dot}{\eta}(n)]|\leq 2^{n}.$
@@
 Write $p^{\prime}=\langle s,x_{0}\rangle$.
-There exists $x_{1}\in[x_{0}]^{\omega}\cap\mathcal{U}$ such that the $\mathbb{M}_{\mathcal{U}}$-condition $\langle s,x_{1}\rangle\geq_{\mathbb{M}_{\mathcal{U}}}\langle s,x_{0}\rangle$ has the following property:
-For every $t\in\text{fin}(x_{1})$, there exists $C_{t}\in\text{fin}(\omega)$ such that
+There exists $x_{1}\in[x_{0}]^{\omega}\cap\mathcal{U}$ such that the $\mathbb{M}_{\mathcal{U}}$-condition $\langle s,x_{1}\rangle\leq_{\mathbb{M}_{\mathcal{U}}}\langle s,x_{0}\rangle$ has the following property:
+For every $t\in[x_{1}]^{<\omega}$, there exists $C_{t}\in[\omega]^{<\omega}$ such that
 $\langle s\cup t,x_{1}\setminus(\max t)^{+}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}}\undertilde{C}\cap(\max t)^{+}=C_{t}.$
@@
 Assume $\langle s,x_{1}\rangle$ is as in the conclusion of the previous lemma. There exists $x_{2}\in[x_{1}]^{\omega}\cap\mathcal{U}$ such that $\langle s,x_{2}\rangle$ has the following property:
-For every $t\in\text{fin}(x_{2})$, every $m\in x_{2}\setminus\max t$ and all $n,n^{\prime}\in x_{2}\setminus m^{+}$, it holds that $C_{t\cup\{n\}}\cap m^{+}=C_{t\cup\{n^{\prime}\}}\cap m^{+}$.
+For every $t\in[x_{2}]^{<\omega}$, every $m\in x_{2}\setminus\max t$ and all $n,n^{\prime}\in x_{2}\setminus m^{+}$, it holds that $C_{t\cup\{n\}}\cap m^{+}=C_{t\cup\{n^{\prime}\}}\cap m^{+}$.
 Proof.
@@
 Proof.
-Note that the conclusion of Lemmas 2.4 and 2.5 also holds for each $\langle s^{\prime},x^{\prime}\rangle\geq_{\mathbb{M}_{\mathcal{U}}}\langle s,x_{2}\rangle$.
+Note that the conclusion of Lemmata 2.4 and 2.5 also holds for each $\langle s^{\prime},x^{\prime}\rangle\leq_{\mathbb{M}_{\mathcal{U}}}\langle s,x_{2}\rangle$.
 Hence, we simply trim $x_{2}$ such that the enumeration of $s\cup x_{2}$ dominates $2^{j+1}$ above $|s|$ and replace $s$ with $s\cup\{\min x_{2}\}$ and $x_{2}$ with $x_{2}\setminus\{\min x_{2}\}$.
@@
 Next, let $N$ be a countable elementary submodel of some large enough $\mathcal{H}_{\chi}$ such that $\{\mathcal{U},\mathbb{M}_{\mathcal{U}},\undertilde{C},\langle s,x_{2}\rangle\}\in N$.
-By induction, construct a sequence $N_{0}\subseteq N_{1}\subseteq...$ of finite subsets of $N$ such that
+By induction, construct a sequence $N_{0}\subseteq N_{1}\subseteq$...
+of finite subsets of $N$ such that
   • (i)
@@
   • (iv)
-    $\forall i\in\omega:\forall t\in\text{fin}(\omega):t\in N_{i}\iff t\subseteq N_{i}$,
+    $\forall i\in\omega:\forall t\in[\omega]^{<\omega}:t\in N_{i}\iff t\subseteq N_{i}$,
   • (v)
-    If $\langle m,l,D\rangle\in(\omega\times\omega\times\text{fin}(\omega))\cap N_{i}$, then $m,l,D\in N_{i}$ (and hence $D\subseteq N_{i}$ by the previous condition).
+    If $\langle m,l,D\rangle\in(\omega\times\omega\times[\omega]^{<\omega})\cap N_{i}$, then $m,l,D\in N_{i}$ (and hence $D\subseteq N_{i}$ by the previous condition).
   • (vi)
@@
 Proof.
-Assume $\langle s\cup t,x^{\prime}\rangle\geq_{\mathbb{M}_{\mathcal{U}}}\langle s,x_{2}\rangle$, $a\in\omega\setminus(\max s)^{+}$ and $i\in\omega\setminus\{0,1\}$ are such that
+Assume $\langle s\cup t,x^{\prime}\rangle\leq_{\mathbb{M}_{\mathcal{U}}}\langle s,x_{2}\rangle$, $a\in\omega\setminus(\max s)^{+}$ and $i\in\omega\setminus\{0,1\}$ are such that
 $\langle s\cup t,x^{\prime}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}}a\in\undertilde{C}\setminus(\max s)^{+}\cap[k_{i-1},k_{i}).$
@@
 If $l^{*}\in[k_{i-1},k_{i})$, we are done, since this means that $\langle s\cup t,x^{\prime}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}}l^{*}\in\text{range}(\underaccent{\dot}{\eta})\cap[k_{i-1},k_{i})$.
-Hence, assume $l^{*}\notin[k_{i-1},k_{i})$, which means that $l^{*}\notin N_{i}$, since $l^{*}$ is certainly not in $N_{i-1}$ (if it were, $a$ would be as well by (iii)).
+Hence, assume $l^{*}\notin[k_{i-1},k_{i})$, i.e., $l^{*}\notin N_{i}$.
 Note that $l^{*}$ witnesses that
 $N\models\exists l:l=\min(x_{2}\setminus a).$
-Hence, by (v), we have that $l^{*}\in N_{i+1}$ and thus $l^{*}\in[k_{i},k_{i+1})$.
+Hence, by (vi), we have that $l^{*}\in N_{i+1}$ and thus $l^{*}\in[k_{i},k_{i+1})$.
 Case II.
@@
 We will show that $i^{*}=i-1$.
-Let $j\in\omega$ be such that $l^{*}$ is the $j$’th elements of $s\cup t$ in increasing order. By Lemmas 2.4 and 2.5, there is $C_{t^{\prime}\cup\{l^{*}\}}\subseteq(l^{*})^{+}$ such that
+Let $j\in\omega$ be such that $l^{*}$ is the $j$’th elements of $s\cup t$ in increasing order. By Lemmata 2.4 and 2.5, there is $C_{t^{\prime}\cup\{l^{*}\}}\subseteq(l^{*})^{+}$ such that
 $\langle s\cup t^{\prime}\cup\{l^{*}\},x_{2}\setminus(l^{*})^{+}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}}\undertilde{C}\cap(l^{*})^{+}=C_{t^{\prime}\cup\{l^{*}\}}.$
 Set $D^{*}:=C_{t^{\prime}\cup\{l^{*}\}}\cap(l_{0},m^{*})$. Since
-$\langle s\cup t^{\prime}\cup\{l^{*}\},x_{2}\setminus(l^{*})^{+}\rangle\leq_{\mathbb{M}_{\mathcal{U}}}\langle s\cup t,x^{\prime}\rangle,$
+$\langle s\cup t^{\prime}\cup\{l^{*}\},x_{2}\setminus(l^{*})^{+}\rangle\geq_{\mathbb{M}_{\mathcal{U}}}\langle s\cup t,x^{\prime}\rangle,$
 and since $l_{0}<a<m^{*}$ by assumption, we must have $a\in D^{*}$. Furthermore, note that $D^{*}\subseteq C_{t^{\prime}\cup\{l^{*}\}}\cap(l_{0},l^{*}]$ and thus $|D^{*}|=:\gamma\leq 2^{j}$.
@@
 $N\models\exists\langle m,l,D\rangle:\begin{cases}m,l\in x_{2}\setminus{l_{0}}^{+},m<l,\text{ and}\\ D\subseteq(l_{0},m),\text{ and }\\ |D|=\gamma,\text{ and }\\ \langle s\cup t^{\prime}\cup\{l\},x_{2}\setminus l^{+}\rangle\Vdash_{\mathbb{M}_{\mathcal{U}}}\undertilde{C}\cap(l_{0},m)=D.\end{cases}$
-Since $l_{0}$ is the $(j-1)$’th element of $s\cup t^{\prime}$, we have $l_{0}>2^{j}$ by Fact 2.6.666Note that the additional requirement in Fact 2.6 that $\max s$ is already larger than $2^{|s|}$ is needed here, since $l_{0}$ could be $\max s$.
+Since $l_{0}$ is the (j1)(j-1)’th element of $s\cup t^{\prime}$, we have $l_{0}>2^{j}$ by Fact 2.6.888Note that the additional requirement in Fact 2.6 that $\max s$ is already larger than $2^{|s|}$ is needed here, since $l_{0}$ could be $\max s$.
 Hence, since $l_{0}\in N_{i^{*}}$, it follows that $\gamma\in N_{i^{*}}$.
@@
 Note that the proof of this claim will finish the proof of the Lemma, since $l^{\dagger}\in N_{i^{*}+1}$ by (v) and thus $a\in N_{i^{*}+1}\setminus N_{i^{*}}$.
-Proof. Assume by contradiction that $l^{\dagger}<a$, i.e.,
+Proof of Claim. Assume by contradiction that $l^{\dagger}<a$, i.e.,
 $l_{0}<m^{\dagger}<l^{\dagger}<a<m^{*}<l^{*}.$
@@
 However, both $D^{\dagger}$ and $D^{*}$ have size $\gamma$ and thus $D^{*}\subseteq(l_{0},m^{\dagger})$, which is a contradiction to the fact that $a\in D^{*}$ and $a>m^{\dagger}$.
-$\blacksquare$   $\dashv$
+This completes the proof of the Claim and Lemma 2.7 as well.
+$\dashv$
 We now only need one final lemma to finish the proof of the proposition and thus of the Main Theorem.
@@
 Let $f$ be an order-preserving bijection from $v_{0}$ to $u_{0}$, extended to a permutation of $\omega$.
-Thus, for each $i\in\omega$, $f$ sends the element selected by $v_{0}$ in $[k_{i},k_{i+1})$ to the element selected by $u_{0}$ in $[k_{i},k_{i+1})$.
+Thus, for each $i\in\omega$, $f$ sends the element selected by $v_{0}$ in $k_{i},k_{i+1})$ to the element selected by $u_{0}$ in $k_{i},k_{i+1})$.
 Since $E$ and $\mathcal{U}$ are non-isomorphic, there exist $v_{1}\in[v_{0}]^{\omega}\cap E$ and $u_{1}\in[u_{0}]^{\omega}\cap\mathcal{U}$ such that $u_{1}\cap f[v_{1}]=\emptyset$.
@@
 Let $G\ast H$ be any $\mathbb{M}_{\mathcal{U}}\ast\undertilde{R}-$generic filter containing $\langle\langle s,x_{2}\rangle,\undertilde{q}^{\prime}\rangle$.
-By Lemma 2.7, we have that in $\mathbf{V}[G\ast H]$, whenever $\text{range}(\undertilde{a}[G\ast H])\setminus(\max s)^{+}$ intersects one of the intervals $[k_{i},k_{i+1})$, then the Mathias real $\eta$ intersects $[k_{i},k_{i+1})$ or one of the adjacent intervals $[k_{i-1},k_{i})$ or $[k_{i+1},k_{i+2})$.
+By Lemma 2.7, we have that in $\mathbf{V}[G\ast H]$, whenever $\text{range}(\undertilde{a}[G\ast H])\setminus(\max s)^{+}$ intersects one of the intervals $k_{i},k_{i+1})$, then the Mathias real $\eta$ intersects $k_{i},k_{i+1})$ or one of the adjacent intervals $k_{i-1},k_{i})$ or $k_{i+1},k_{i+2})$.
 Since $\text{range}(\eta)$ is almost contained in $u$, the same is true for $u$ in place of $\eta$ above some $n\geq(\max s)^{+}$.
 Hence, $\text{range}(\undertilde{a}[G\ast H])\setminus n$ is disjoint from $v$.
-$\dashv$   $\dashv$
+$\dashv$